# Discovery and Analysis of Afterglows from Poorly Localised GRBs with the Gravitational-wave Optical Transient Observer (GOTO) All-sky Survey

### v1

@@ Abstract
Gamma-ray bursts (GRBs), particularly those detected by wide-field instruments such as the Fermi/GBM, pose a challenge for optical follow-up due to their large initial localisation regions, leaving many GRBs without identified afterglows. The Gravitational-wave Optical Transient Observer (GOTO), with its wide field of view, dual-site coverage, and robotic rapid-response capability, bridges this gap by rapidly identifying and localising afterglows from alerts issued by space-based facilities including Fermi, SVOM, Swift, and the EP, providing early optical positions for coordinated multi-wavelength follow-up. In this paper, we present optical afterglow localisation and multi-band follow-up of seven Fermi/GBM and MAXI/GSC triggered long GRBs (240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B) discovered by GOTO in 2024. Spectroscopy for six GRBs (no spectroscopic data for GRB 241002B) with VLT/X-shooter and GTC/OSIRIS yields precise redshifts spanning $z\approx 0.40$–3.16 and absorption-line diagnostics of host and intervening systems. Radio detections for four events confirm the presence of long-lived synchrotron emission. Prompt-emission analysis with Fermi and MAXI data reveals a spectrally hard population, with two bursts lying $3\sigma$ above the Amati relation. Although their optical afterglows resemble those of typical long GRBs, the prompt spectra are consistently harder than the long-GRB average. Consistent modelling of six GOTO-discovered GRB afterglows yields jet half-opening angles of a few degrees and beaming-corrected kinetic energies $E_{\rm jet}\sim 10^{51}$$10^{52}$ erg, consistent with the canonical long-GRB population. These findings suggest that optical discovery of poorly localised GRBs may be subject to observational biases favouring luminous events with high spectral peak energy ($E_{\rm p}$), while also providing insight into jet microphysics and central engine diversity.
@@ 1 Introduction
The study of gamma-ray bursts (GRBs) has advanced significantly since their discovery in the 1960s (Klebesadel et al., 1973; Strong et al., 1974), driven by a combination of dedicated space-based surveys and ground-based follow-up observations. From confirming their cosmological origins (Meegan et al., 1992; Costa et al., 1997; Metzger et al., 1997; van Paradijs et al., 1997) to uncovering possible progenitors for long and short GRBs (Mazets et al., 1981; Kouveliotou et al., 1993; Woosley & Bloom, 2006; Zhang et al., 2012; Abbott et al., 2017b), these high-energy events have now been recognised as one of the most powerful explosions in the universe (Mészáros, 2013; Kumar & Zhang, 2015; Levan et al., 2016; LHAASO Collaboration et al., 2023). Observations across the electromagnetic spectrum have not only enhanced our understanding of GRB mechanisms but also established their ability to probe the distant universe (Fiore, 2001; Tanvir et al., 2009; Petitjean & Vergani, 2011; Saccardi et al., 2023, 2025) and constrain cosmological parameters (Amati & Della Valle, 2013; Demianski et al., 2017; Luongo & Muccino, 2021; Moresco et al., 2022), marking them as invaluable tools in modern astrophysics.
@@
Early optical and multi-wavelength follow-up has revealed a broad diversity in afterglow behaviours, including evidence for reverse shocks (Zhang & Kobayashi, 2005; Mundell et al., 2007; Mimica et al., 2009, 2010; Laskar et al., 2013; Yi et al., 2020), energy injection episodes (Björnsson et al., 2004; Zhang et al., 2006; Laskar et al., 2015), and structured jets (Lamb & Kobayashi, 2017; Beniamini et al., 2020; Oganesyan et al., 2020). Combined with X-ray and radio data, optical observations enable comprehensive modelling of afterglows, shedding light on the energetics and structure of GRB jets (Margutti et al., 2013). However, a persistent challenge in GRB afterglow detection and follow-up arises from the poor initial localisation provided by wide-field gamma-ray monitors such as Fermi Gamma-ray Burst Monitor (GBM; Meegan et al. 2009). With the highest GRB detection rate and strong sensitivity to prompt gamma-ray emission, Fermi/GBM enables detailed temporal and spectral studies (Meegan et al., 2009; von Kienlin et al., 2020), but typically provides localisation uncertainties spanning several square degrees. In contrast, missions like Neil Gehrels Swift Observatory (Swift hereafter; Gehrels et al. 2004), Einstein Probe (EP; Yuan et al. 2015, 2022), and Space-based multi-band astronomical Variable Objects Monitor (SVOM; Wei et al. 2016) offer arcsecond–arcminute level localisations but detect comparatively fewer bursts. Recovering counterparts to these poorly localised GRBs provides a valuable opportunity to expand our understanding of GRB diversity and reduce selection biases. However, the large error regions often exceed the field of view (FoV) of conventional optical telescopes, complicating timely afterglow identification. Additionally, the intrinsic faintness of some afterglows (Liang et al., 2007; Dereli et al., 2017) and circumburst extinction (Savaglio & Fall, 2004; Schulze et al., 2011) can further hinder follow-up. Without alternative localisation strategies, a significant fraction of GRBs, particularly those detected by Fermi/GBM, remain uncharacterised, limiting our ability to probe jet physics, energetics, and progenitor properties.
@@ 2.1 GOTO Overview
Each site hosts two independent mounts, each carrying eight UTs with a combined instantaneous FoV of $\approx 44\,deg^{2}$. Across the 16 UTs at both sites, GOTO achieves a total FoV of $\approx 88\,deg^{2}$, making it well-suited to search for optical counterparts of poorly localised transients such as gravitational-wave events (e.g., Gompertz et al., 2020), GRBs detected by facilities such as Fermi, SVOM and EP (e.g., Mong et al., 2021; Belkin et al., 2024), as well as other fast and exotic transients including rapidly-evolving supernovae and tidal disruption events. The telescopes are equipped with ON Semiconductor KAF-50100 CCDs, which provide broad sensitivity across the optical range, with the deployed Baader filters setting the effective bandpass. In survey mode, a wide $L$-band (400700400-700 nm) encompassing the Sloan gri filters is used, providing sensitivity to a wide range of transients and maximising discovery potential.
@@ 2.3 Identifying the GRB Optical Counterpart
    Temporal behaviour: evidence of fading between successive GOTO epochs, or a later non-detection deeper than the initial detection; where possible, the decline should be consistent with power-law afterglow behaviour.
@@ 2.4 Observational Performance and Discovery Statistics
GOTO’s GRB follow-up programme has developed from its prototype stage into a fully operational dual-site facility (Dyer et al., 2024), delivering a series of notable discoveries. The prototype system (GOTO-4), comprising four UTs at the La Palma site, achieved first light in June 2017 and was officially inaugurated in July 2017, initiating routine operations (Steeghs et al., 2022). During this early phase, GOTO secured its first GRB optical afterglow detection with GRB 171205A, associated with SN 2017iuk (Steeghs et al., 2017; Izzo et al., 2019). In the initial 3 years between June 2017 and June 2020, the GOTO-4 system responded to 77 Fermi/GBM and 29 Swift/BAT triggers (see Mong et al., 2021; Steeghs et al., 2022), demonstrating the scientific potential of the facility even in its prototype configuration.
With full deployment at both sites, as of 31 December 2024 (with GRBs thereafter observed to be included in a future study), GOTO has conducted follow-up observations for over 257 Fermi -, 43 Swift -, 28 EP-, and 7 GECAM-triggered events. No targeted observations were conducted for SVOM events during this period, as GOTO began following up SVOM triggers in 2025, coinciding with the scheduled start of SVOM’s science operations in February 2025. On average, GOTO’s first targeted observation latency was approximately 11.2 hr, with response times ranging from 270 s–69.77 hr post-trigger. The average latency for the last observation in each follow-up series was 26.36 hr, while the alert latency averaged 4.8 hr, underscoring the challenges in achieving timely and efficient follow-ups.
@@
GRB 230911A was the first LGRB for which GOTO discovered an optical afterglow (Belkin et al., 2023); details are published in Belkin et al. (2024). After this first case, in 2024, GOTO discovered optical afterglows of 2 MAXI-triggered (GRB 240122A and GRB 240225B) serendipitously and 5 Fermi-triggered (GRBs 240619A, 240910A, 240916A, 241002B and 241228B) LGRBs in responsive mode, which are studied in detail in the present work. The details of these 7 GRBs in our sample are discussed in the following section. In addition to these LGRBs, within 2024 itself, GOTO also identified the optical afterglow of the SGRB 241105A (Julakanti et al., 2024b), which was localised by Fermi/GBM with an uncertainty of $\sim 4\,\mathrm{deg}$ (Fermi GBM Team, 2024e). GOTO rapidly responded to the trigger, tiling 277.9 square degrees within the 90% GBM localisation region and covering $\sim 84\%$ of the total probability within 1.6 hr. The afterglow was discovered at $L\sim 17.2$ mag, later confirmed through multi-wavelength follow-up and spectroscopy to lie at a redshift of $z=2.681$ (Izzo et al., 2024). Although this event is not part of the core LGRB sample analysed here, it highlights GOTO’s capabilities to detect optical afterglows from both long and short GRBs, even under challenging localisation conditions. A detailed analysis of GRB 241105A is presented in a separate paper by Dimple et al. (2025).
@@ 3.2 Localisation Coverage and Optical Afterglow Discoveries
Figure 3 provides an overview of the sky localisation and follow-up coverage for the GRBs analysed in this study (GRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B), based on GOTO observations obtained within the first 24 h post-trigger. The dark and light grey contours represent the $1\sigma$ and $2\sigma$ localisation regions extracted from the HEALPix (Górski et al., 2005) probability skymaps. Light blue shading denotes the GOTO fields observed within 10 hr of each burst, demonstrating the system’s wide-area and rapid-response capabilities. Cyan stars mark the locations of confirmed optical afterglows, showing that GOTO’s coverage either enclosed or closely bordered the true source positions in all cases. In addition, the finding charts highlighting the confirmed afterglows, derived from GOTO observations, are presented in panels (a)–(g) of Figure 4. For context, each GOTO cutout is paired with a corresponding image from the DESI Legacy Survey (LS) DR10 covering the same $\sim 6.3^{\prime}\times 6.3^{\prime}$ FoV; for GRB 240122A, a Pan-STARRS DR1 image is shown instead.
@@ 3.2.1 MAXI/GSC GRBs 240122A and 240225B: arcminute-scale localisation
GOTO does not follow up MAXI triggers in responsive mode, as MAXI alerts are not distributed in a machine-readable format that can be ingested automatically by the sentinel. Instead, these fields are only covered serendipitously in survey mode. In Figure 3, the top row shows the two events detected by MAXI/GSC, GRB 240122A and GRB 240225B, each displayed within a $3^{\circ}\times 3^{\circ}$ FoV. The red shaded regions indicate the approximate localisation areas from MAXI/GSC reports. For GRB 240122A, the afterglow is well centred within the localisation and fully encompassed by the GOTO field. In contrast, for GRB 240225B, the GOTO tiling intersected the elongated error region, providing timely coverage that included the eventual afterglow position.
@@
GOTO-S serendipitously observed the localisation region of GRB 240122A during its routine all-sky survey on 2024-01-22, discovering the optical afterglow (GOTO24eu/AT2024apy) at J2000 coordinates RA = 06h12m12s.91 and Dec = -190838′′.81. The afterglow was detected at 11:11:43 UT (T0+43.68 min) with an $L$-band magnitude of $17.06\pm 0.04$ mag (Kumar et al., 2024b), see also Table 1 for details. As shown in Figure 3 (top-left), the GOTO FoV comfortably covers the compact MAXI/GSC localisation, placing the afterglow well within the imaged area. This case demonstrates GOTO’s ability to capture transient counterparts during its high-cadence survey mode, even without a targeted trigger.
@@
Figure 4: Finding charts of GRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B in the GOTO $L$-band (400–700 nm) observed by GOTO. Each cutout is a $300\times 300$ pixel region centred on the transient, corresponding to a FoV of $\sim 6.3\arcmin\times 6.3\arcmin$ at the GOTO pixel scale of $1.26\arcsec$/pix. For comparison, survey images from the DESI Legacy Survey DR10 are shown (except for GRB 240122A, where a Pan-STARRS DR1 image is used), matched to the same FoV. Details of each object are listed in Table 1.
@@ 3.2.2 Fermi/GBM GRBs 240619A, 240910A, 240916A, 241002B, and 241228B: degree-scale localisation
Fermi/GBM triggers are distributed in real time via machine-readable GCN notices, which the GOTO sentinel ingests automatically. This enables fully responsive follow-up, with observations scheduled immediately after the alert is received. Follow-up observations are carried out as soon as observing conditions and visibility constraints permit. The five different panels in Figure 3 correspond to events detected by Fermi/GBM GRBs 240619A, 240910A, 240916A, 241002B, and 241228B, each displayed over a wider $20^{\circ}\times 20^{\circ}$ field. A quantitative summary of GOTO follow-up coverage for the Fermi/GBM GRBs in our sample is provided in Table 2. For each event, we list the GBM localisation uncertainty (expressed as the circularised 90% containment radius in degrees), number of images obtained to cover the 90% GBM localisation region, the total sky area imaged, the fraction of localisation probability enclosed within the observed fields and the mean $5\sigma$ limiting magnitude. The two MAXI/GSC GRBs are not included here, as their compact localisations were fully covered serendipitously by single $\sim 9$ deg2 survey pointings.
@@
GOTO-N initiated follow-up observations of GRB 241228B on 2024-12-28 at 04:26:19 UT (T0+0.22 h), continuing through to 23:24:22 UT (T0+19.19 h). As depicted in Figure 3 (bottom-left panel), the GOTO tiling efficiently covered the high-probability localisation region, including the afterglow position. The afterglow (GOTO24jmz/AT2024afgu) was identified by GOTO-N at J2000 coordinates RA = 08h31m05s.46 and Dec = +065054′′.07, with an initial detection at 04:32:24 UT (T0+19.3 min) at an $L$-band magnitude of $14.54\pm 0.01$ mag. Multiple detections followed throughout the observing sequence, with the final GOTO-N detection recorded at 13:00:42 UT (T0+8.79 h) at $19.70\pm 0.10$ mag (Kumar et al., 2024d). Afterglow candidate for GRB 241228B lies on the 94.5% probability contour, formally outside the GBM 90% localisation region. While most GRB afterglows are found within the 90% contour, a small fraction are expected to lie just beyond it, making GRB 241228B a noteworthy case.

### v2

@@ Abstract
Gamma-ray bursts (GRBs), particularly those detected by wide-field instruments such as the Fermi/GBM, pose challenges for optical follow-up because of their large initial localisation regions, leaving many GRBs without identified afterglows. The Gravitational-wave Optical Transient Observer (GOTO), with its wide field of view, dual-site coverage, and robotic rapid-response capability, bridges this gap by rapidly identifying and localising afterglows from alerts issued by space-based facilities including Fermi, SVOM, Swift, and the EP, providing early optical positions for coordinated multi-wavelength follow-up. In this paper, we present optical afterglow localisation and multi-band follow-up of five Fermi/GBM (240619A, 240910A, 240916A, 241002B, and 241228B) and two MAXI/GSC (240122A and 240225B) triggered long GRBs discovered by GOTO in 2024. Spectroscopy for six GRBs (no spectroscopy for GRB 241002B) with VLT/X-shooter and GTC/OSIRIS yields precise redshifts spanning $z\approx 0.40$–3.16 and absorption-line diagnostics of hosts and intervening systems. Radio detections for four events (240122A, 240619A, 240910A, and 240916A) confirm the presence of long-lived synchrotron emission. Prompt-emission analysis with Fermi and MAXI data reveals a spectrally hard population, with two bursts lying $3\sigma$ above the Amati relation. Although their optical afterglows resemble those of typical long GRBs, the prompt spectra are consistently harder than the long-GRB average. Broadband afterglow modelling of six GOTO-discovered GRBs yields jet half-opening angles of a few degrees and beaming-corrected kinetic energies $E_{\rm jet}\sim 10^{51}$$10^{52}$ erg, consistent with the canonical long-GRB population. These findings suggest that optical discovery of poorly localised GRBs is likely subject to observational biases favouring luminous events with high spectral peak energy ($E_{\rm p}$), while also providing insight into jet microphysics and central engine diversity.
@@ 1 Introduction
The study of gamma-ray bursts (GRBs) has advanced significantly since their discovery in the 1960s (Klebesadel et al., 1973; Strong et al., 1974), driven by a combination of dedicated space-based surveys and ground-based follow-up observations. From confirming their cosmological origins (Meegan et al., 1992; Costa et al., 1997; Metzger et al., 1997; van Paradijs et al., 1997) to uncovering possible progenitors for long and short GRBs (Mazets et al., 1981; Kouveliotou et al., 1993; Woosley & Bloom, 2006; Zhang et al., 2012; Abbott et al., 2017b), these high-energy events have now been recognised as among the most luminous explosive phenomena in the universe (Mészáros, 2013; Kumar & Zhang, 2015; Levan et al., 2016; LHAASO Collaboration et al., 2023). Observations across the electromagnetic spectrum have not only enhanced our understanding of GRB mechanisms but also established their ability to probe the distant universe (Fiore, 2001; Tanvir et al., 2009; Petitjean & Vergani, 2011; Saccardi et al., 2023, 2025) and constrain cosmological parameters (Amati & Della Valle, 2013; Demianski et al., 2017; Luongo & Muccino, 2021; Moresco et al., 2022), marking them as invaluable tools in modern astrophysics.
@@
Early optical and multi-wavelength follow-up has revealed a broad diversity in afterglow behaviours, including evidence for reverse shocks (Zhang & Kobayashi, 2005; Mundell et al., 2007; Mimica et al., 2009, 2010; Laskar et al., 2013; Yi et al., 2020), energy injection episodes (Björnsson et al., 2004; Zhang et al., 2006; Laskar et al., 2015), and structured jets (Lamb & Kobayashi, 2017; Beniamini et al., 2020; Oganesyan et al., 2020). Combined with X-ray and radio data, optical observations enable comprehensive modelling of afterglows, shedding light on the energetics and structure of GRB jets (Margutti et al., 2013). However, a persistent challenge in GRB afterglow detection and follow-up arises from the poor initial localisation provided by wide-field gamma-ray monitors such as Fermi Gamma-ray Burst Monitor (GBM; Meegan et al. 2009). With the highest GRB detection rate and strong sensitivity to prompt gamma-ray emission, Fermi/GBM enables detailed temporal and spectral studies (Meegan et al., 2009; von Kienlin et al., 2020), but typically provides localisation uncertainties spanning several square degrees. In contrast, missions like Neil Gehrels Swift Observatory (Swift hereafter; Gehrels et al. 2004), Einstein Probe (EP; Yuan et al. 2015, 2022), and Space-based multi-band astronomical Variable Objects Monitor (SVOM; Wei et al. 2016) offer arcsecond- to arcminute-level localisations but detect comparatively fewer bursts. Recovering counterparts to these poorly localised GRBs provides a valuable opportunity to expand our understanding of GRB diversity and reduce selection biases. However, the large error regions often exceed the field of view (FoV) of conventional optical telescopes, complicating timely afterglow identification. Furthermore, the intrinsic faintness of some afterglows (Liang et al., 2007; Dereli et al., 2017) and circumburst extinction (Savaglio & Fall, 2004; Schulze et al., 2011) can further hinder follow-up. Without alternative localisation strategies, a significant fraction of GRBs, particularly those detected by Fermi/GBM, remain uncharacterised, limiting our ability to probe jet physics, energetics, and progenitor properties.
@@ 2.1 GOTO Overview
Each site hosts two independent mounts, with the eight UTs on each mount aligned to form a tiled array with small overlaps, yielding a combined FoV of $\approx 44\,\mathrm{deg}^{2}$ per mount. With two mounts per site, this provides $\approx 88\,\mathrm{deg}^{2}$ of instantaneous coverage, and across both sites the network spans $\approx 176\,\mathrm{deg}^{2}$. This wide coverage makes GOTO particularly well-suited to search for optical counterparts of poorly localised transients such as gravitational-wave events (e.g., Gompertz et al., 2020), GRBs detected by facilities such as Fermi, SVOM and EP (e.g., Mong et al., 2021; Belkin et al., 2024), as well as other fast and exotic transients including rapidly-evolving supernovae and tidal disruption events. The telescopes are equipped with ON Semiconductor KAF-50100 CCDs, which provide broad sensitivity across the optical range, with the deployed Baader filters setting the effective bandpass. In survey mode, a wide $L$-band (400700400-700 nm) encompassing the Sloan gri filters is used, providing sensitivity to a wide range of transients and maximising discovery potential.
@@ 2.3 Identifying the GRB Optical Counterpart
    Temporal behaviour: evidence of fading between successive GOTO epochs, or a later non-detection deeper than the initial detection; where possible, the decline should be consistent with a power-law afterglow behaviour.
@@ 2.4 Observational Performance and Discovery Statistics
The GOTO GRB follow-up programme has developed from its prototype stage to a fully operational dual-site facility (Dyer et al., 2024), delivering a series of notable discoveries. The prototype system (GOTO-4), comprising four UTs at the La Palma site, achieved first light in June 2017 and was officially inaugurated in July 2017, initiating routine operations (Steeghs et al., 2022). During this early phase, GOTO secured its first GRB optical afterglow detection with GRB 171205A, associated with SN 2017iuk (Steeghs et al., 2017; Izzo et al., 2019). In the initial 3 years between June 2017 and June 2020, the GOTO-4 system responded to 77 Fermi/GBM and 29 Swift/BAT triggers (see Mong et al., 2021; Steeghs et al., 2022), demonstrating the scientific potential of the facility even in its prototype configuration.
With full deployment at both sites, as of 31 December 2024 (with GRBs observed thereafter will be included in a future study), GOTO has conducted follow-up observations for over 257 Fermi -, 43 Swift -, 28 EP-, and 7 GECAM-triggered events. No targeted observations were conducted for SVOM events during this period, as GOTO began following up SVOM triggers in 2025, coinciding with the scheduled start of SVOM’s science operations in February 2025. On average, GOTO’s first targeted observation latency was approximately 11.2 hr, with response times ranging from 270 s–69.77 hr post-trigger. The average latency for the last observation in each follow-up series was 26.36 hr, while the alert latency averaged 4.8 hr, underscoring the challenges in achieving timely and efficient follow-ups.
@@
GRB 230911A was the first LGRB for which GOTO discovered an optical afterglow (Belkin et al., 2023); details are published in Belkin et al. (2024). After this first case, in 2024, GOTO discovered optical afterglows of 2 MAXI-triggered (GRB 240122A and GRB 240225B) serendipitously and 5 Fermi-triggered (GRBs 240619A, 240910A, 240916A, 241002B and 241228B) LGRBs in responsive mode, which are studied in detail in this work. The details of these 7 GRBs in our sample are discussed in the following section. In addition to these LGRBs, within 2024 itself, GOTO also identified the optical afterglow of the SGRB 241105A (Julakanti et al., 2024b), which was localised by Fermi/GBM with an uncertainty of $\sim 4\,\mathrm{deg}$ (Fermi GBM Team, 2024e). GOTO rapidly responded to the trigger, tiling 277.9 square degrees within the 90% GBM localisation region and covering $\sim 84\%$ of the total probability within 1.6 hr. The afterglow was discovered at $L\sim 17.2$ mag, later confirmed through multi-wavelength follow-up and spectroscopy to lie at a redshift of $z=2.681$ (Izzo et al., 2024). Although this event is not part of the core LGRB sample analysed here, it highlights GOTO’s capabilities to detect optical afterglows from both long and short GRBs, even under challenging localisation conditions. A detailed analysis of GRB 241105A is presented in a separate paper by Dimple et al. (2025).
@@ 3.2 Localisation Coverage and Optical Afterglow Discoveries
Figure 3 provides an overview of the sky localisation and follow-up coverage for the GRBs analysed in this study (GRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B), based on GOTO observations obtained within the first 24 h post-trigger. The dark and light grey contours represent the $1\sigma$ and $2\sigma$ localisation regions extracted from the HEALPix (Górski et al., 2005) probability skymaps. Light blue shading denotes the GOTO fields observed within 10 hr of each burst, demonstrating the system’s wide-area and rapid-response capabilities. Cyan stars mark the locations of confirmed optical afterglows, showing that GOTO’s coverage either enclosed or closely bordered the true source positions in all cases. In addition, the finding charts highlighting the confirmed afterglows, derived from GOTO observations, are presented in panels (a)–(g) of Figure 4. For context, each GOTO cutout is paired with a corresponding image from Legacy Survey (LS) DR10 covering the same $\sim 6.3^{\prime}\times 6.3^{\prime}$ FoV; for GRB 240122A, a Pan-STARRS DR1 image is shown instead.
@@ 3.2.1 MAXI/GSC GRBs 240122A and 240225B: arcminute-scale localisation
GOTO does not follow up MAXI triggers in responsive mode, as MAXI alerts are not distributed in a machine-readable format that can be ingested automatically by the sentinel. Instead, these fields are only covered serendipitously in survey mode. In Figure 3, the top row shows the two events detected by MAXI/GSC, GRB 240122A and GRB 240225B, each displayed within a $3\degree\times 3\degree$ FoV. The red shaded regions indicate the approximate localisation areas from MAXI/GSC reports. For GRB 240122A, the afterglow is well centred within the localisation and fully encompassed by the GOTO field. In contrast, for GRB 240225B, the GOTO tiling intersected the elongated error region, providing timely coverage that included the eventual afterglow position.
@@
GOTO-S serendipitously observed the localisation region of GRB 240122A during its routine all-sky survey on 2024-01-22, discovering the optical afterglow (GOTO24eu/AT2024apy) at J2000 coordinates RA = 06h12m12s.91 and Dec = -190838′′.81. The afterglow was detected at 11:11:43 UT (T0+43.68 min) with an $L$-band magnitude of $17.58\pm 0.04$ mag (Kumar et al., 2024b), see also Table 1 for details. As shown in Figure 3 (top-left), the GOTO FoV comfortably covers the compact MAXI/GSC localisation, placing the afterglow well within the imaged area. This case demonstrates the ability of GOTO to capture transient counterparts during its high-cadence survey mode, even without a targeted trigger.
@@
Figure 4: Finding charts of GRBs 240122A, 240225B, 240619A, 240910A, 240916A, 241002B, and 241228B in the GOTO $L$-band (400–700 nm) observed by GOTO. Each cutout is a $300\times 300$ pixel region centred on the transient, corresponding to a FoV of $\sim 6.3\arcmin\times 6.3\arcmin$ at the GOTO pixel scale of $1.26\arcsec$/pix. For comparison, survey images from the Legacy Survey DR10 are shown (except for GRB 240122A, where a Pan-STARRS DR1 image is used), matched to the same FoV. Details of each object are listed in Table 1.
@@ 3.2.2 Fermi/GBM GRBs 240619A, 240910A, 240916A, 241002B, and 241228B: degree-scale localisation
Fermi/GBM triggers are distributed in real time via machine-readable GCN notices, which the GOTO sentinel ingests automatically. This enables fully responsive follow-up, with observations scheduled immediately after the alert is received. Follow-up observations are carried out as soon as observing conditions and visibility constraints permit. The five different panels in Figure 3 correspond to events detected by Fermi/GBM GRBs 240619A, 240910A, 240916A, 241002B, and 241228B, each displayed over a wider $20\degree\times 20\degree$ field. A quantitative summary of GOTO follow-up coverage for the Fermi/GBM GRBs in our sample is provided in Table 2. For each event, we list the GBM localisation uncertainty (expressed as the circularised 90% containment radius in degrees), number of images obtained to cover the 90% GBM localisation region, the total sky area imaged, the fraction of localisation probability enclosed within the observed fields and the mean $5\sigma$ limiting magnitude. The two MAXI/GSC GRBs are not included here, as their compact localisations were fully covered serendipitously by single $\sim 9$ deg2 survey pointings.
@@
GOTO-N initiated follow-up observations of GRB 241228B on 2024-12-28 at 04:26:19 UT (T0+0.22 h), continuing through to 23:24:22 UT (T0+19.19 h). As depicted in Figure 3 (bottom-left panel), the GOTO tiling efficiently covered the high-probability localisation region, including the afterglow position. The afterglow (GOTO24jmz/AT2024afgu) was identified by GOTO-N at J2000 coordinates RA = 08h31m05s.46 and Dec = +065054′′.07, with an initial detection at 04:32:24 UT (T0+0.32 h) at an $L$-band magnitude of $14.54\pm 0.01$ mag. Multiple detections followed throughout the observing sequence, with the final GOTO-N detection recorded at 13:00:42 UT (T0+8.79 h) at $19.70\pm 0.10$ mag (Kumar et al., 2024d). The afterglow candidate for GRB 241228B falls on the 94.5% probability contour, formally outside the GBM 90% localisation region. While most GRB afterglows are found within the 90% contour, a small fraction are expected to lie just beyond it, making GRB 241228B a noteworthy case.